Calculus II, Fall 2014 (http://www.math.nthu.edu.tw/˜ wangwc/)

Homework Assignment for Week 09

1. Section 14.5: Problems 9, 15, 19, 25, 29, 35.

2. Section 14.6: Problems 1, 9, 13, 19, 25, 33, 39, 45, 59, 63, 66, 67.